OpenStudy (anonymous):

solve : 3x/(2) + 1/(x)= 3x/(4)

OpenStudy (anonymous):

3x/2 - 3x/4 = -1/x => 3x/4 = -1/x => x^2 = -4/3 => x = (-4/3)^1/2

OpenStudy (anonymous):

why is -4/3 to the 1/2?

OpenStudy (anonymous):

what is why? is something wrong? do u know? if x^y = z then x = z^(1/y)

OpenStudy (anonymous):

why do you put -4/3 to the 1/2?

OpenStudy (anonymous):

http://www.twiddla.com/495681 use this link......

OpenStudy (anonymous):

whats the password?

OpenStudy (anonymous):

password.....